[Brahms-l] Shift report 20040101 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Thu Jan 01 2004 - 23:56:12 EST
------------------------------------------------------------
Date: Thu Jan  1 23:56:11 2004, Shift: 16:00-24:00
Supervisor: Steve Sanders, with: Oana Ristea
Log book#15 pages 20 - 20

Shift Summary:
------------------------------------------------------------
1. Beam from 16-20 and from 23 until now.

2. Run 9307 - D1, D2, D3, D4 = 1/4B, D5 = 350B - 146k (tr1);
   run 9309 - D1, D2, D3, D4 = 1/4B, D5 = 350B - 302k (tr1);
   run 9310 - D1, D2, D3, D4 = 1/4B, D5 = 350B - 44k (tr1);
   run 9311 - D1, D2, D3, D4 = 1/4B, D5 = 350B.

3. Checked NIM-ECL converter for ZDC gate.  Looks good.
Also checked that all LVL0 gates lead to ZDC gate.  Checks.
Why all the zero channels?
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Thu Jan 1 23:56:21 2004

This archive was generated by hypermail 2.1.8 : Thu Jan 01 2004 - 23:56:38 EST